# On Geometry and Topology of Hessian Manifolds

### v4

@@ 5.3 Compact Orientable Hessian 3-folds
and denote $\Omega:=\{(x,y,t)\in\mathbb{R}^{3}:Q(x,y,t)>0\}$. Also, define the smooth function $f\in C^{\infty}(\Omega)$ by
@@
for every $(\gamma,n)\in G$ and every $(\tau,\rho)\in\mathbb{H}^{2}\times\mathbb{R}$. Therefore the 1-form potential $df$ of $((\Omega,\nabla),\nabla(df))$ is invariant under the $G$-action, and thus descends to the quotient $M_{\varphi}$. In particular, the product Riemannian metric $ds^{2}$ on $\mathbb{H}^{2}\times\mathbb{R}$ descends to a Riemannian metric $h$ on $M_{\varphi}$.
@@
In particular, we arrived at the following theorem:
@@
Let $\Sigma_{g}$ be a closed orientable surface of genus $g>1$, and let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. Then there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a closed and orientable radiant Hessian $3$-manifold of Koszul type with negative scalar curvature.
@@
The examples constructed above are consistent with the results established earlier. As predicted by Theorem 2.46, for a closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Proposition 2.7 and Proposition 2.19.
@@
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is the closed orientable surface of genus $g\geq 1$.
@@
In the special case that $(M,h)$ is a flat Riemannian manifold, the desired theorem is proved by the classification of space groups.
From now on, we assume that $(M,h)$ is not flat. Then by Theorem 2.48, there exists a $\Sigma_{g}$-bundle $f\colon M\rightarrow\mathbb{S}^{1}$ for some closed orientable surface $\Sigma_{g}$ of genus $g$.
Since $M$ is aspherical by Corollary 2.14, we obtain that $\Sigma_{g}$ ditto. Therefore, in particular, $\Sigma_{g}$ is not homeomorphic to $\mathbb{S}^{2}$, and hence $g\geq 1$.
In particular, we have $\pi_{1}(M)=\pi_{1}(\Sigma_{g})\rtimes_{\varphi}\mathbb{Z}$ where $\varphi$ is the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$. Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then there exists a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$. Denote by $\hat{\varphi}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the lifting of $\varphi\colon\Sigma_{g}\rightarrow\Sigma_{g}$ to the universal covering spaces. Since $\pi_{1}(M)$ acts on $(\mathbb{R}^{n}\times\mathbb{R},g\oplus dt^{2})$ by isometry, in particular the smooth mapping
is an isometry. Therefore we conclude that $\hat{\varphi}$ and hence $\varphi$ are isometries.
Denote $i\colon\Sigma_{g}\rightarrow M$ the inclusion map of a fiber of the fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$. Consider the subgroup $A$ of the isometry group $G$ of $(\Sigma_{g},i^{*}h)$ generated by $\varphi$, and denote by $\bar{A}$ its closure in $G$. Then it is readily seen that $\bar{A}$ is a compact abelian subgroup of $G$, and hence the identity component of $\bar{A}$ is a torus, Therefore, $\varphi$ is isotopic in $G$ to a group element of finite order.
Quod erat demonstrandum. ∎
@@
Theorem 5.15 is also consistent with earlier discoveries: Consider a closed orientable surface $\Sigma_{g}$ of genus $g$. Recall that pseudo-Anosov maps and Dehn twists are of infinite order. By Nielsen–Thurston classification, the mapping torus $M_{\varphi}$ of a pseudo-Anosov mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ admits a hyperbolic metric, which is forbidden by our previous Theorem 4.10. For the very same reason, we obtain that the nilpotent Heisenberg manifold cannot support any Hessian metric.
We close this section with a final remark: It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two. Therefore, in particular, the examples presented in this section are different to those constructed using Example 1.7 and Theorem 1.8.
5.4 Topology of Hessian 4-Manifolds
In this section, we list several properties of Hessian manifolds in dimension four.
@@ 5.4 Topology of Hessian 4-Manifolds
Corollary 5.18.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is null-cobordant, i.e. there exist an orientable smooth 5-manifold $N$ with bundary $\partial N=M$.
@@
See [32] for a proof of the fact that an oriented closed smooth 4-manifold is null-cobordant if and only if it has signature zero. ∎
Recall that all closed orientable 4-manifolds have $\operatorname{spin}^{c}$ structures, and therefore their third integral Stiefel-Whitney classes vanish. The following property is specific for Hessian 4-manifolds.
@@
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. If $M$ is spin then it admits an almost complex structure.
@@
Recall that $M$ is spin if its second Stiefel–Whitney class $w_{2}(TM)\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$ vanishes.
Moreover, by Ehresmann-Wu theorem, $M$ admits an almost complex structure if and only if there exists $c\in\mathrm{H}^{2}(M;\mathbb{Z})$ such that
$3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=c^{2}$ (3)
and $w_{2}(TM)$ is the $\bmod 2$ reduction of $c$. For our case, by Lemma 5.17, $\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=0$. If $w_{2}(TM)=0$, then $c=0$ satisfies equation( 3) and its $\bmod 2$ reduction is trivially equal to $w_{2}(TM)$. ∎
Furthermore, the following proposition shows that there are only two possibilities for the intersection form of an oriented closed Hessian 4-manifold, depending on the existence of $\operatorname{spin}$ structure.
@@
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold with the second Betti number $b_{2}(M)=2m$. If $M$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
if $M$ is not spin, then $Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$.
@@
Recall that $M$ is spin if its second Stiefel–Whitney class $w_{2}:=w_{2}(TM)$ vanishes. By Wu’s formula, we have $Sq^{2}(a)=a\smile a=a\smile w_{2}$ for all $a\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$. It is then readily seen that $Q_{M}$ is even if and only if $w_{2}=0$. The desired result then follows from Hasse-Minkowski classification theorem of indefinite unimodular symmetric bilinear form over $\mathbb{Z}$, and the fact that $\operatorname{sign}(Q_{M})=0$. ∎
To sum up, we suggest several Riemannian geometric properties of Hessian 4-manifolds, showing that a Hessian metric puts severe restrictions on other Riemannian metrics that a $4$-manifold can possess.
Proposition 5.21.
Let $((M,\nabla),h)$ be a compact Hessian 4-manifold. Then for any Riemannian metric $g$ on $M$, the Riemannian 4-manifold $(M,g)$ is not of negative sectional curvature.

### v6

@@ 5.2 Closed Orientable Hessian 3-Manifolds
and denote $\Omega:=\{(x,y,t)\in\mathbb{R}^{3}:Q(x,y,t)>0\}$. Also, define a smooth function $f\in C^{\infty}(\Omega)$ by
@@
for every $(\gamma,n)\in G$ and every $(\tau,\rho)\in\mathbb{H}^{2}\times\mathbb{R}$. Therefore the 1-form potential $df$ of $((\Omega,\nabla),\nabla df)$ is invariant under the $G$-action, and thus descends to the quotient $M_{\varphi}$. In particular, the product Riemannian metric $ds^{2}$ on $\mathbb{H}^{2}\times\mathbb{R}$ descends to a Riemannian metric $h$ on $M_{\varphi}$.
@@
We thus arrived at the following theorem:
@@
Let $\Sigma_{g}$ be a closed orientable surface of genus $g>1$, and let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. Then there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a closed and orientable Hessian $3$-manifold of Koszul type.
@@
The examples constructed above are consistent with the results established earlier. As predicted by Theorem 4.1, for a closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Corollary 2.6 and Proposition 2.16.
@@
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is a closed orientable surface of genus $g\geq 1$.
@@
By Theorem 4.15, without loss of generality, we may assume that $(M,\nabla)$ is the product of a compact hyperbolic affine manifold with a flat torus of dimension $d\in\{0,\dots,3\}$.
If $d\geq 2$, then the desired result holds trivially.
If $d=1$, then $M$ is homeomorphic to the $3$-torus $\mathbb{T}^{3}$ by Theorem 5.6.
Now, assume that $d=0$. Then $(M,\nabla)$ itself is a compact orientable hyperbolic affine manifold. The desired result, in this case, follows directly from Corollary 4.12. ∎
Recall that a Dehn twist on a closed orientable surface $\Sigma_{g}$ of genus $g>0$ is of infinite order. Therefore, in particular, Theorem 5.9 implies that the nilpotent Heisenberg manifold
$M:=H_{3}(\mathbb{R})/H_{3}(\mathbb{Z})$
cannot support any Hessian metric.
@@
5.3 Closed Orientable Hessian 4-Manifolds
In this final section, we classify all closed orientable Hessian $4$-manifolds topologically.
@@ 5.3 Closed Orientable Hessian 4-Manifolds
Recall that all closed orientable 4-manifolds have $\operatorname{spin}^{c}$ structures. The following property is specific for Hessian 4-manifolds.
Proposition 5.12.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. If $(M,h)$ is spin then $M$ admits an almost complex structure with vanishing first Chern class. Moreover, the second Chern classes of all almost complex structures on $M$ are zero.
@@
Recall that $(M,h)$ is spin if the second Stiefel–Whitney class $w_{2}(TM)\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$ of $M$ vanishes.
Moreover, by Ehresmann-Wu theorem, $M$ admits an almost complex structure $J$ with first Chern class $c\in\mathrm{H}^{2}(M;\mathbb{Z})$ if and only if the following equation
$3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=c^{2}$ (3)
is satisfied and $w_{2}(TM)$ is the $\bmod 2$ reduction of $c$. For our case, by Lemma 5.11, $3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=0+0=0$. If $w_{2}(TM)=0$, then $c=0$ satisfies equation (3) and its $\bmod 2$ reduction is trivially equal to $w_{2}(TM)$.
Now, equip the tangent bundle $TM\rightarrow M$ with an almost complex structure $J$. Then, by the theory of characteristic classes, the first Pontryagin class can be expressed as
$p_{1}(TM)=c_{1}(TM)^{2}-2c_{2}(TM)$
in terms of Chern classes. Since $p_{1}(TM)=0$ by Proposition 1.5, and $c_{1}(TM)^{2}=0+0=0$ again by Ehresmann-Wu theorem, we conclude that $c_{2}(TM)=0$. ∎
Furthermore, the following result shows that there are only two possibilities for the intersection form of an oriented closed Hessian 4-manifold, depending on the existence of $\operatorname{spin}$ structure.
@@
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold with the second Betti number $b_{2}(M)=2m$. If $(M,h)$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
if $(M,h)$ is not spin, then $Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$.
@@
Recall that $(M,h)$ is spin if the second Stiefel–Whitney class $w_{2}:=w_{2}(TM)$ of $M$ vanishes. By Wu’s formula, we have $Sq^{2}(a)=a\smile a=a\smile w_{2}$ for all $a\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$. It is then readily seen that $Q_{M}$ is even if and only if $w_{2}=0$. The desired result then follows from Hasse-Minkowski classification theorem of indefinite unimodular symmetric bilinear form over $\mathbb{Z}$, and the fact that $\operatorname{sign}(Q_{M})=0$. ∎
In fact, a closed orientable Hessian 4-manifold is either spin or Riemannian flat:
@@
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Suppose that the Riemannian manifold $(M,h)$ is not flat. Then $(M,h)$ is spin.
@@
By Proposition 4.2, there exist a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ with connected orientable fiber $N$. Denote by $dt\in\Omega^{1}(\mathbb{S}^{1})$ the volume form of the unit circle, and denote by $\phi]\in\operatorname{Mod}(N)$ the monodromy of $f\colon M\rightarrow\mathbb{S}^{1}$. Then, by Wang sequence, the second cohomology group $\mathrm{H}^{2}(M;\mathbb{Z})$ decomposes as
$\mathrm{H}^{2}(M;\mathbb{Z})=\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}\oplus A$
where
$A:=\{[\alpha]\smile f^{*}[dt]:[\alpha]\in\mathrm{H}^{1}(N;\mathbb{Z})\}\cong\mathrm{H}^{1}(N;\mathbb{Z})/\operatorname{Im}(\phi^{*}-\operatorname{Id})$
and $\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}$ is the submodule of $\mathrm{H}^{2}(N;\mathbb{Z})$ consists of $\phi$-invariant classes. It is readily seen that for any $\beta]\in\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}$ or $\beta]\in A$, we have $\beta]\smile[\beta]=0$. Also, by equivariant Poincaré duality, the intersection form $Q_{M}$ restricts to a non-degenerated pairing between $\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}$ and $A$. Therefore, in particular, we have $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]$
and hence $(M,h)$ is spin by Proposition 5.13. ∎
Finally, with all preparations done, we now state and prove our main theorem:
Theorem 5.15.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    a Bieberbach 4-manifold;
  2. 2.
    the mapping torus of an isometry of a hyperbolic 3-manifold;
  3. 3.
    a $\Sigma_{g}$-bundle over $\mathbb{T}^{2}$ with periodic monodromy,
where $\Sigma_{g}$ here is the closed orientable surface of genus $g>1$.